Casey Nelson is a divisional manager for Pigeon Company. His annual pay raises are largely determined by his division’s return on investment (ROI), which has been above 24% each of the last three years. Casey is considering a capital budgeting project that would require a $4,800,000 investment in equipment with a useful life of five years and no salvage value. Pigeon Company’s discount rate is 20%. The project would provide net operating income each year for five years as follows:

Sales $ 4,500,000

Variable expenses 2,040,000

Contribution margin 2,460,000

Fixed expenses: Advertising, salaries, and other fixed out-of-pocket costs $ 810,000

Depreciation 960,000

Total fixed expenses 1,770,000

Net operating income $ 690,000

Required: 1. What is the project’s net present value? 2. What is the project’s internal rate of return to the nearest whole percent? 3. What is the project’s simple rate of return? 4-a. Would the company want Casey to pursue this investment opportunity? 4-b. Would Casey be inclined to pursue this investment opportunity?
Answer #1
Net operating income 690000
Add: Depreciation 960000
Net cash flows 1650000
1
Now Year 1 Year 2 Year 3 Year 4 Year 5
Investment cost -4800000
Net cash flows 1650000 1650000 1650000 1650000 1650000
Total cash flows -4800000 1650000 1650000 1650000 1650000 1650000
PV factor @ 20% 1 0.833 0.694 0.579 0.482 0.402
Present value of cash flows -4800000 1374450 1145100 955350 795300 663300
Net present value 133500
2
PV factor internal rate of return = 4800000/1650000 = 2.909
The PV factor 2.909 for 5 years is closest to 21%
Internal rate of return = 21%
3
Simple rate of return = Net operating income/Investment cost
Simple rate of return = 690000/4800000 = 14.4%
4a
Yes, the company would want Casey to pursue this investment as Net Present value is positive
4b
No, Casey would not be inclined to pursue this investment as as his ROI will decrease